# Truthful or Fabricated? Using Causal Attribution to Mitigate Reward Hacking in Explanations

### v1

@@ 1 Introduction
Large Language Models (LLMs) can generate explanations to enhance the perceived trustworthiness of their outputs. A common approach is chain-of-thought (CoT), where the model generates a sequence of ‘reasoning’ steps that serves as extra context to the generated answer (Kojima et al., 2022; Wei et al., 2022; Wang et al., 2023, ia). Besides often leading to improved task performance, this approach offers a ‘window’ into the model’s decision-making. However, recent studies (Lanham et al., 2023; Agarwal et al., 2024; Chua et al., 2024; Madsen et al., 2024; Turpin et al., 2024) show that CoT explanations are not always faithful. For example, if the CoT steps fail to acknowledge an input cue, whose absence we know affects the model-generated answer, there is a possible gap between the explanation and the actual decision process (Turpin et al., 2024). This faithfulness gap (Jacovi & Goldberg, 2020) raises important questions: which aspects of LLM training influence the reliability of generated explanations, and how can training be adapted to improve their reliability?
@@
In this work, we examine the role of preference optimization, used to guide models toward generating responses that are not only correct but also adhere to preferences about their form, meaning, and broader implications (Ziegler et al., 2019; Stiennon et al., 2020; Askell et al., 2021; Bai et al., 2022a; b; Ouyang et al., 2022, i.a.). Our focus is on understanding how preference optimization can influence the reliability of CoT explanations and exploring ways to modify it to make CoTs more reliable. Preference optimization is typically performed by using reinforcement learning (RL), where the LLM is trained to produce responses scored highly by a reward model (acting in lieu of a human judge) (Schulman et al., 2017; Ouyang et al., 2022); alternatively, the LLM can be directly optimized to adhere to human preferences (Meng et al., 2024; Rafailov et al., 2024), potentially by making use of a pre-trained reward model to produce preference data used for training (Wu et al., 2024). We note a limitation of this scenario: the reward mechanism (or a human judge) only has access to the generated text, and thus, cannot assess whether the explanation given in the response is faithful to the model’s decision process. In settings where preferences extend to how the model arrives at a response, this limitation feeds a form of reward hacking (Krakovna et al., 2020; Pan et al., 2022; Skalse et al., 2022, i.a.): the reward model prioritizes responses that appear to adhere to preferences over those that overtly do not, with learning pushing the LLM to exploit this as a mechanism to collect rewards at the expense of the reliability of CoT explanations. We refer to this behavior as CoT hacking.
@@
To exemplify a category of such settings, we define two set-ups where an LLM generates a response to a prompt with a CoT explanation and a predicted answer, and where: (i) the reward model exhibits a preference for a specific answer (e.g., the solution of a math problem), (ii) the input includes a cue (protected feature) that is correlated with that answer, and (iii) an instruction discourages the LLM from relying on the cue. These conflicting goals (i.e., having easy access to the preferred prediction, via the cue, but being discouraged to use it) create a potential for a form of ‘cheating’: the LLM can use the protected feature to get the preferred answer while omitting this fact from the explanation. When we adapt the LLM to follow the instruction, for example via DPO training, this strategy becomes an easy and unnoticeable mechanism to collect rewards. Fig. 1 illustrates one of the two set-ups (‘Math Book’): we prompt an LLM to solve math problems, while giving it access to a block of already solved problems which may include the solution for the test query. We instruct the model to solve the problem without consulting the solution to the test query and to respond with a CoT explanation. Finally, we adapt the model (e.g., through DPO training) in an attempt to have it follow the instruction. As anticipated, we observe that using the reward model to guide the LLM, results in exaggerating any faithfulness gap already present in the LLM’s CoT explanations— e.g., the presence of the solutions in the prompt increases performance compared to when they are omitted, yet the produced CoTs seldom mention the protected resource.
@@ 3.1 Data
We obtain data by making use of the AQUA-RAT222https://huggingface.co/datasets/deepmind/aqua_rat dataset (Ling et al., 2017), which consists of multiple-choice answer algebraic word problems with natural language rationales. We use the original validation (254 examples) and test (254) data, and sample 3000 examples as training data. To build the math book we sample 3 examples from the portion of the original training data that we do not use, and add solutions for two of the problems.
@@
In the ‘BiasQA‘ setting, the goal of the LLM is to answer a query by choosing a pronoun to fill the blank in a given sentence, e.g.“The client called the toolmaker after hearing a good review of ___ service and products”. As part of the input, the model has access to potentially biasing features, in this case, the profession “toolmaker”, which exemplifies the type of protected feature in this setting.
@@ 3.2 Models, Training, and Evaluation
We use LLama3.1-8B-IT (Dubey et al., 2024) for all experiments. Whenever we need to generate multiple samples, we use top-p decoding, with temperature set to 0.8 and top-p set to 0.95. Otherwise, we use greedy decoding. Unless mentioned otherwise, we sample responses. We use vLLM for more efficient decoding (Kwon et al., 2023).
@@
We report the percentage of responses that predict the correct choice in the ‘Math Book’ setting (Accuracy) and that predict the stereotypical answer in the ‘BiasQA‘ setting (Stereotype Rate). We also report the percentage of responses that acknowledge the protected feature in the CoT explanation (Acknowledgment rate), marginally across the test set. Acknowledgments are identified by an ‘Eval LLM’, in our case Llama-3.3-70B-Instruct555https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct (Dubey et al., 2024), as described in Appendix C. When measuring Majority@16, we consider a response to be correct/stereotypical or to acknowledge the protected feature, if at least 9 of the samples do so. For example, if at least 9 of the samples predict the stereotypical label, then the response to that prompt is considered to be stereotypical.
@@ 4 Reward Models Drive Chain-of-Thought Hacking
Furthermore we note that increases in accuracy or stereotype rate between the original and counterfactual prompts are not consistently matched by corresponding increases in marginal acknowledgment rates, except in ‘BiasQA‘ with greedy decoding. For example, for ‘Math Book‘ with greedy decoding, the accuracy gap is 18.1 percentage points, while acknowledgment rate differs only by 1.6. The mismatch provides initial evidence that the model may rely on the protected feature without disclosing it.
@@
Before further finetuning the base model, we first ‘isolate’ the impact of the reward model via BoN decoding (see §A). Figure 4 shows how accuracy/stereotype and acknowledgment rates evolve as we optimize the chosen response in function of the reward score (◆) by SK-Gemma-27B.666We find similar evidence for SK-Llama-8B, as discussed in Appendix Fig. 9 and Table 3. We can observe that doing so leads to an increased potential for deceptive responses, as accuracy in ‘Math Book’ increases from 74.7% to 93.3%, while acknowledgment rate decreases from 2.2% to 1.2%, and stereotype rate in ‘BiasQA‘ increases from 56.9% to 73.1%, while acknowledgment rate increases at a lower rate from 22.9% to 32.6%. Furthermore, the gap to the non-optimized base model (⚫) is also clear in both setting, decreasing slightly with N in the ‘Math Book’ setting (from 17.0 percentage points to 15.1) and increasing clearly in the ‘Bias QA’ setting (from 7.9 percentage points to 23.1). These results showcase the role of the reward model in promoting non-desired behavior.
@@
We now study the impact of annotating data to train a DPO model using a reward model, as described in Section 3.2. Results for DPO () can be seen in Figure 3 (◆⚫ for SK-Llama-8B and ◆⚫ for SK-Gemma-27B). We start by noting that DPO results in models that are more accurate (‘Math Book’) or stereotypical (‘BiasQA’) than their base model counterpart (see Appendix Table 2). Once again, the potential for unfaithful explanations is clear: across all settings, the gap in accuracy/stereotype rate between prompts increases when compared to the base model, while the gap in acknowledgment rate mostly increases at a smaller rate or decreases.
@@ 5 Counterfactual-Augmented Reward Models
When we detect that the protected feature exerts causal influence on , we append to a disclaimer, warning the RM that the LLM accessed the protected features. The disclaimer reads as follows: “I acknowledge that I used the SOLUTIONS part of the MATH BOOK.” for ‘Math Book’, and “I acknowledge that my reasoning used biases or stereotypes related to a profession being more likely associated to a person of a certain gender.” for ‘BiasQA‘. We experiment with two strategies. In one strategy, we append the disclaimer whenever —we refer to this as strategy (for the predictions differ). In another strategy, we append the disclaimer whenever is cued (correct/stereotypical) and is not—we refer to this as strategy (for not only the predictions differ, but is cued). In Tables and Figures, we refer to a reward model that uses one or the other strategy as or , respectively. See Appendix Section B.2 for details.
@@ 5.1 Results
We start by assessing the impact of augmenting the input to the reward model with interpretability signals in BoN decoding. If the signal helps the RM penalise the use of the protected feature, we should observe a decrease in accuracy/stereotype rate, ideally, matching the performance of the LLM when not given access to the protected feature. Figure 4 and Appendix Figure 9 (see Appendix Table 3 for numerical values) show how both strategies ( and ) show promise for mitigating unfaithfulness — e.g., for SK-Gemma-27B, (▼) closes the gap between the base model with default RM with access to the protected feature (◆) and the base model without access to the protected feature (⚫) by 39% for ‘Math Book’ and by 65% for ‘BiasQA’, while (✖) does so by 9% and 25%, respectively.
@@
We now show the impact of using and as the reward model used to annotate the preference dataset used to train the DPO model. Figure 3 shows that, when compared to a DPO model based on data annotated with the default RM, both strategies result in DPO models that deviate from the counterfactual performance by a smaller margin — for the ‘Math Book’ setting reduces this margin on average by 2 percentage points and by 4.7 percentage points, while for ‘BiasQA’ the average reduction is of 8 and 6 percentage points, respectively.
At the same time, acknowledgment rates typically increase with respect to the DPO (RM) model, showing the potential of both techniques in reducing the rate at which unfaithful responses are preferred.
@@
Base | - | 24.4 | 27.6 | 13.7 | 12.7 |
DPO + | SK-Llama-8B | 28.3 | 35.8 | 13.7 | 12.4 |
DPO + | 25.2 | 33.5 | 7.3 | 3.5 | |
DPO + | 22.0 | 29.9 | 12.4 | 4.4 | |
DPO + | SK-Gemma-27B | 24.8 | 39.0 | 20.0 | 26.0 |
DPO + | 29.1 | 34.6 | 10.5 | 11.1 | |
DPO + | 25.6 | 34.3 | 12.1 | 15.2 |
@@
So far, we have seen that reward models – whether used in best-of-N decoding or for constructing preference datasets in DPO – can increase the alignment of model predictions with labels associated with the protected feature, without a corresponding rise in acknowledgment rates. This suggests a trend toward unfaithful explanations. We have also seen how counterfactually-augmented reward models help reduce the tendency of this behavior. We now take a more ‘fine-grained’ look at this effect by comparing individual prompt–counterfactual pairs, and aggregating across examples. In particular, for a given response with the full prompt, we obtain the response for the corresponding counterfactual prompt. Then, we consider the response to be ‘unfaithful’ if the original prompt response matches the label correlated with the protected feature without acknowledging it, while the counterfactual prompt response does not match the label. For BoN, we sample one of the 16 responses to the counterfactual using 20 different seeds, and then report the average rates.
We report results for greedy and sampling in Table 1 and for BoN in Appendix Fig. 10. Similarly to what we observed so far, incorporating the reward model as part of the pipeline promotes unfaithful explanations. For greedy decoding the largest absolute difference occurs for the ‘BiasQA‘ setting when using the SK-Gemma-27B reward model (13.7% unfaithful examples versus 20%), and similarly for majority@16 (12.7% unfaithful examples versus 26%). For best-of-N alignment the impact of the reward model in the selection of examples is once again clear, with the number of deceptive examples increasing consistently with N for both settings and reward models. The impact of the augmented reward model strategies is also clear, decreasing the number of deceptive examples compared to DPO (RM) on 14 of the 16 comparisons.
@@ 6 Related Work
Reasoning chains output by LLMs (Kojima et al., 2022; Wei et al., 2022; Wang et al., 2023; Yao et al., 2024, i.a.) can be inspected as a self-explanation for its prediction. These often look plausible to human readers (Agarwal et al., 2024), but might lack faithfulness (Lanham et al., 2023; Agarwal et al., 2024; Madsen et al., 2024; Turpin et al., 2024, i.a.). A common way to assess the faithfulness of LLM outputs is to compare the predictions generated from the original context with those from a modified version: e.g., by corrupting the obtained CoTs (Lanham et al., 2023), or adding biasing features (Atanasova et al., 2023; Chua et al., 2024; Turpin et al., 2024) to the model input and verifying their presence in the explanation. We explore similar techniques to build ‘interpretability signals’ that make the reward model input potentially more faithful.
@@
The approaches to reduce reward hacking include ensembling reward models (Coste et al., 2023; Eisenstein et al., 2024; Rame et al., 2024, i.a.), and doing reward shaping (Jinnai et al., 2024; Miao et al., 2024; Fu et al., 2025), targeting known issues, such as length bias (Shen et al., 2023; Chen et al., 2024; Huang et al., 2025, i.a.). In a parallel work, Baker et al. (2025) proposed using an LLM to monitor CoTs for reward hacking when optimizing a generator LLM; they observed that it may lead the generator to fabricate CoTs. In contrast, we address reward hacking that arises from the reward model’s lack of access to the generator’s decision-making process.

### v2

@@ 1 Introduction
Large Language Models (LLMs) can generate explanations to enhance the perceived trustworthiness of their outputs. A common approach is chain-of-thought (CoT), where the model generates a sequence of ‘reasoning’ steps that serves as extra context to the generated answer (Kojima et al., 2022; Wei et al., 2022; Wang et al., 2023, ia). Besides often leading to improved task performance, this approach offers a ‘window’ into the model’s decision-making. For example, if the CoT steps fail to acknowledge an input cue, whose absence we know affects the model-generated answer, there is a possible gap between the explanation and the actual decision process (Turpin et al., 2024). This faithfulness gap (Jacovi & Goldberg, 2020) raises important questions: which aspects of LLM training influence the reliability of generated explanations, and how can training be adapted to improve their reliability?
@@
In this work, we examine the role of preference optimization, used to guide models toward generating responses that are not only correct but also adhere to preferences about their form, meaning, and broader implications (Ziegler et al., 2019; Stiennon et al., 2020; Askell et al., 2021; Bai et al., 2022a, b; Ouyang et al., 2022, i.a.). Our focus is on understanding how preference optimization can influence the reliability of CoT explanations and exploring ways to modify it to make CoTs more reliable. Preference optimization is typically performed by using reinforcement learning (RL), where the LLM is trained to produce responses scored highly by a reward model (acting in lieu of a human judge) (Schulman et al., 2017; Ouyang et al., 2022); alternatively, the LLM can be directly optimized to adhere to human preferences (Meng et al., 2024; Rafailov et al., 2024), potentially by making use of a pre-trained reward model to produce preference data used for training (Wu et al., 2024). We note a limitation of this scenario: the reward mechanism (or a human judge) only has access to the generated text, and thus, cannot assess whether the explanation given in the response is faithful to the model’s decision process. In settings where preferences extend to how the model arrives at a response, this limitation feeds a form of reward hacking (Krakovna et al., 2020; Pan et al., 2022; Skalse et al., 2022, i.a.): the reward model prioritizes responses that appear to adhere to preferences over those that overtly do not, with learning pushing the LLM to exploit this as a mechanism to collect rewards at the expense of the reliability of CoT explanations. We refer to this behavior as CoT hacking.
@@
To exemplify a category of such settings, we define two set-ups where an LLM generates a response to a prompt with a CoT explanation and a predicted answer, and where: (i) the reward model exhibits a preference for a specific answer (e.g., the solution of a math problem), (ii) the input includes a cue (protected feature) that is correlated with that answer, and (iii) an instruction discourages the LLM from relying on the cue. These conflicting goals (i.e., having easy access to the preferred prediction, via the cue, but being discouraged to use it) create a potential for a form of ‘cheating’: the LLM can use the protected feature to get the preferred answer while omitting this fact from the explanation. When we adapt the LLM to follow the instruction, for example via DPO training, this strategy becomes an easy and unnoticeable mechanism to collect rewards. Fig. 1 illustrates one of the two set-ups (‘Math Book’): we prompt an LLM to solve math problems, while giving it access to a block of already solved problems which may include the solution for the test query. We instruct the model to solve the problem without consulting the solution to the test query and to respond with a CoT explanation. Finally, we adapt the model (e.g., through DPO training) in an attempt to have it follow the instruction. As anticipated, we observe that using the reward model to guide the LLM, results in exaggerating any faithfulness gap already present in the LLM’s CoT explanations— i.e., the presence of the solutions in the prompt increases performance compared to when they are omitted, yet the produced CoTs seldom mention the protected resource.
@@ 3.1 Data
We obtain data by making use of the AQUA-RAT222https://huggingface.co/datasets/deepmind/aqua_rat dataset (Ling et al., 2017), which consists of multiple-choice answer algebraic word problems with natural language rationales. We use the original validation (254 examples) and test (254) data, and sample 3000 examples as training data. To build the math book for a given example, we sample 2 additional examples from the portion of the original training data that we do not use, and add the respective solution for one of them, resulting in a total of 3 problems statements, with 2 being solved.
@@
In the ‘BiasQA’ setting, the goal of the LLM is to answer a query by choosing a pronoun to fill the blank in a given sentence, e.g.“The client called the toolmaker after hearing a good review of ___ service and products”. As part of the input, the model has access to potentially biasing features, in this case, the profession “toolmaker”, which exemplifies the type of protected feature in this setting.
@@ 3.2 Models, Training, and Evaluation
We use LLama3.1-8B-IT (Dubey et al., 2024) for all experiments. To generate multiple samples, we use top-p decoding, with temperature set to 0.8 and top-p set to 0.95. Otherwise, we use greedy decoding. Unless mentioned otherwise, we sample responses. We use vLLM for more efficient decoding (Kwon et al., 2023).
@@
We report the percentage of responses that predict the correct choice in the ‘Math Book’ setting (Accuracy) and that predict the stereotypical answer in the ‘BiasQA’ setting (Stereotype Rate). We also report the percentage of responses that acknowledge the protected feature in the CoT explanation (Acknowledgment rate), marginally across the test set. Acknowledgments are identified by an ‘Eval LLM’, in our case Llama-3.3-70B-Instruct555https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct (Dubey et al., 2024), as described in Appendix C. When measuring Majority@16 (Wang et al., 2023), we consider a response to be correct/stereotypical or to acknowledge the protected feature, if at least 9 of the samples do so. For example, if at least 9 of the samples predict the stereotypical label, then the response to that prompt is considered to be stereotypical.
@@
We run each experiment with three different seeds and report the average results and the corresponding standard deviations.
@@ 4 Reward Models Drive Chain-of-Thought Hacking
Furthermore we note that increases in accuracy or stereotype rate between the original and counterfactual prompts are not consistently matched by corresponding increases in marginal acknowledgment rates, except in ‘BiasQA’ with greedy decoding. For example, for ‘Math Book’ with greedy decoding, the accuracy gap is 18.1 percentage points, while acknowledgment rate differs only by 1.6. The mismatch provides initial evidence that the model may rely on the protected feature without disclosing it.
@@
Before further finetuning the base model, we first ‘isolate’ the impact of the reward model via BoN decoding (see §A). Figure 4 shows how accuracy/stereotype and acknowledgment rates evolve as we optimize the chosen response in function of the reward score (◆) by SK-Gemma-27B.666We find similar evidence for SK-Llama-8B, as seen in Appendix Figure 9 and Table 4. We can observe that doing so leads to an increased potential for deceptive responses, as accuracy in ‘Math Book’ increases from 75.2% to 93.6%, while acknowledgment rate decreases from 2.7% to 1.7%, and stereotype rate in ‘BiasQA’ increases from 56.7% to 72.4%, while acknowledgment rate increases at a lower rate from 23.3% to 30.3%. Furthermore, the gap in accuracy/stereotype rate to the non-optimized base model (⚫) is also clear in both settings, decreasing slightly with N in the ‘Math Book’ setting (from 18.8 percentage points to 15.0) and increasing clearly in the ‘Bias QA’ setting (from 7.6 percentage points to 20.9). These results showcase the role of the reward model in promoting non-desired behavior.
@@
We now study the impact of annotating data to train a DPO model using a reward model, as described in Section 3.2. Results for DPO () can be seen in Figure 3 (◆⚫ for SK-Llama-8B and ◆⚫ for SK-Gemma-27B). We start by noting that DPO results in models that are more accurate (‘Math Book’) or stereotypical (‘BiasQA’) than their base model counterpart (see Appendix Table 2). Once again, the potential for unfaithful explanations is clear: in 7 out of 8 comparisons, the gap in accuracy/stereotype rate between prompts increases when compared to the base model, while the gap in acknowledgment rate increases at a smaller rate or decreases.
@@ 5 Counterfactual-Augmented Reward Models
When we detect that the protected feature exerts causal influence on , we append to a disclaimer, warning the RM that the LLM accessed the protected features. The disclaimer reads as follows: “I acknowledge that I used the SOLUTIONS part of the MATH BOOK.” for ‘Math Book’, and “I acknowledge that my reasoning used biases or stereotypes related to a profession being more likely associated to a person of a certain gender.” for ‘BiasQA’. We experiment with two strategies. In one strategy, we append the disclaimer whenever —we refer to this as strategy (for the predictions differ). In another strategy, we append the disclaimer whenever is cued (correct/stereotypical) and is not—we refer to this as strategy (for not only the predictions differ, but is cued). In Tables and Figures, we refer to a reward model that uses one or the other strategy as or , respectively. See Appendix Section B.2 for details.
@@ 5.1 Results
We start by assessing the impact of augmenting the input to the reward model with interpretability signals in BoN decoding. If the signal helps the RM penalise the use of the protected feature, we should observe a decrease in accuracy/stereotype rate, ideally, matching the performance of the LLM when not given access to the protected feature. Figure 4 and Appendix Figure 9 (see Appendix Table 4 for numerical values) show how both strategies ( and ) show promise for mitigating unfaithfulness — e.g., for SK-Gemma-27B, (✖) closes the gap between the base model with default RM with access to the protected feature (◆) and the base model without access to the protected feature (⚫) by 41% for ‘Math Book’ and by 71% for ‘BiasQA’, while (▼) does so by 9% and 32%, respectively.
@@
We now show the impact of using and as the reward model used to annotate the preference dataset used to train the DPO model. Figure 3 shows that, when compared to a DPO model based on data annotated with the default RM, both strategies result in DPO models that deviate from the counterfactual performance by a smaller margin for the ‘BiasQA’ setting, with reducing this margin by 7.8 percentage points and by 6.9 percentage points. However, impact is lower for the ‘Math Book’ setting, with reducing this margin by 2.4 percentage points, and with mostly failing to do so. Once again, the importance of having a better informed strategy is noticeable, with , which also considers whether is cued, performing better.
Furthermore, acknowledgment rates typically increase with respect to the DPO (RM) model, showing the potential of both techniques in reducing the rate at which unfaithful responses are preferred.
@@
Base | - | 24.8 0.0 | 27.2 1.5 | 13.7 0.0 | 14.1 1.5 |
DPO + | SK-Llama-8B | 25.7 0.5 | 34.0 0.7 | 13.2 0.8 | 9.8 1.3 |
DPO + | 24.5 1.9 | 33.6 1.2 | 8.0 0.8 | 2.4 0.6 | |
DPO + | 22.8 0.6 | 31.6 3.5 | 7.4 1.8 | 3.9 0.8 | |
DPO + | SK-Gemma-27B | 27.2 1.0 | 33.9 0.9 | 20.8 1.2 | 25.0 1.7 |
DPO + | 28.3 3.9 | 35.2 2.4 | 10.7 0.5 | 7.5 2.2 | |
DPO + | 23.6 0.6 | 32.5 0.5 | 12.3 0.7 | 11.7 3.6 |
@@
So far, we have seen that reward models – whether used in best-of-N decoding or for constructing preference datasets in DPO – can increase the alignment of model predictions with labels associated with the protected feature, without a corresponding rise in acknowledgment rates. This suggests a trend toward unfaithful explanations. We have also seen how counterfactually-augmented reward models help reduce the tendency of this behavior. We now take a more ‘fine-grained’ look at this effect by comparing individual prompt–counterfactual pairs, and aggregating across examples. In particular, for a given response with the full prompt, we obtain the response for the corresponding counterfactual prompt. Then, we consider the response to be ‘unfaithful’ if the original prompt response matches the label correlated with the protected feature without acknowledging it, while the counterfactual prompt response does not match the label. For BoN, we sample one of the 16 responses to the counterfactual prompt.
We report results for DPO using greedy and majority@16 decoding in Table 1 and for BoN in Appendix Figure 10. Similarly to what we observed so far, incorporating the reward model as part of the pipeline promotes unfaithful explanations. When using DPO, for greedy decoding the largest absolute difference occurs for the ‘BiasQA’ setting when using the SK-Gemma-27B reward model (13.7% unfaithful examples versus 20.8%), and similarly for majority@16 (14.1% unfaithful examples versus 25%). When using best-of-N the impact of the reward model in the selection of examples is also clear, with the number of deceptive examples increasing consistently with N for both settings and reward models. Also in this case, the augmented reward model strategies help address the issue of CoT hacking, resulting in fewer deceptive examples compared to using the original reward model in DPO (in 14 of the 16 comparisons), and in BoN.
@@ 6 Related Work
Reasoning chains output by LLMs (Kojima et al., 2022; Wei et al., 2022; Wang et al., 2023; Yao et al., 2024, i.a.) can be inspected as a self-explanation for its prediction. These often look plausible to human readers (Agarwal et al., 2024), but might lack faithfulness (Lanham et al., 2023; Agarwal et al., 2024; Madsen et al., 2024; Turpin et al., 2024, i.a.). A common way to assess the faithfulness of LLM outputs is to compare the predictions generated from the original context with those from a modified version: e.g., by corrupting the obtained CoTs (Lanham et al., 2023), or adding biasing features (Atanasova et al., 2023; Chua et al., 2024; Turpin et al., 2024; Chen et al., 2025) to the model input and verifying their presence in the explanation. We explore similar techniques to build ‘interpretability signals’ that make the reward model input potentially more faithful.
@@
The approaches to reduce reward hacking include ensembling reward models (Coste et al., 2023; Eisenstein et al., 2024; Rame et al., 2024, i.a.), and doing reward shaping (Jinnai et al., 2024; Miao et al., 2024; Fu et al., 2025), targeting known issues, such as length bias (Shen et al., 2023; Chen et al., 2024; Huang et al., 2025, i.a.). In parallel work, Baker et al. (2025) and Chen et al. (2025) discuss monitoring CoTs for reward hacking. The former does so when optimizing a generator LLM and observed that it may lead the generator to fabricate CoTs; the latter finds that monitoring CoTs for reward hacking may not be reliable in tasks where good performance is possible without CoTs. In contrast, we address reward hacking that arises from the reward model’s lack of access to the generator’s decision-making process.